Fri 11th June '10
1 pm T 21 C. Weighed (R) 22 kgs (L) 24 kgs - pretty much the same as last time.
Investigated cross comb, not too bad. Removed 2 small pieces and sewed one of these to a TB and reinserted in between bars, as well as another bar with new comb on it from before. Also put 1 of new TBs with wooden strip in between other bars. Encouraged to do this as the really good news is that they have built some really good combs on the 2 bars we put in 2 weeks ago!!! Removed some brace comb. No sign of queen cells or varroa. Plenty of brood, mostly worker with some drone. Sprayed with homeopathy (7) and small amount of icing sugar.
Roof is leaking and record book got wet. Decided to start this blog instead. Also asked Mick to put roofing felt on both roofs.
Sun 6th June '10
5.30 pm T 17 C. Lots of activity, 40+ bees entering and exiting, many bearing pollen. Good buzz sound. No wax being drawn from new TBs yet, but plenty of bees on them.
Fri 4th June '10
5 pm 2 new TB's with wooden strip rubbed with bees own new wax inserted either end. One spare one prepared and left in hive. Very busy T 20 C.
Thurs 3rd June '10
11 am T 22 C. Lots of activity, pollen bearing bees. Some visible evidence of cross comb. No queen cells seen from looking from each end. Fresh nectar stores at L end. Possible need for more TBs?
Weds 26th May '10
1 pm Removed varroa tray early, after 5 days, as have to do something about cross comb.
Weighed (R) 22.5 kgs (L) 24 kgs. T 18 C.
Investigated cross comb at entrance side and at back! Not as bad as we had feared. Removed cross comb. Sewed some onto a TB and reinserted, a few bars in, at brood end. Also moved one of bars we did before to in between other bars. Bees have also been building out in right direction, either side, especially on one bar. Also tidied up some brace comb. Dusted with icing sugar.
Sue counted 22 varroa on tray. Jim checked and found 33. This when divided by 5 is a daily varroa count of 6.6. This multiplied by 30 (as suggested on our conventional course for this time of year) comes to 198. The worrying level is apparently 1,000, so this is really low - great. Dissected some of the drone cells we had removed in cross comb. Found approx 1 in 5 had a single mite in it.
Fri 21st May '10
7 pm Advised by neighbour of a swarm flying over Linton Farm. Rushed home to check bees. Looked in, either end of bars. Seem fine. Still flying well. Seem content and very strong buzz from main brood area. Unlikely to have swarmed. T 18 C still! However building fresh comb again at R-angle to bars. This time on entrance side.
Fri 21st May '10
2 pm Have attached another varroa mesh to tray, so bees can't get on it. Thought this would prevent any live mites getting back on bees and also bees coming into contact with Vaseline, as used petroleum based product this time (cheaper). Put tray under hive to do a 7 day count. T 23 C. Bees content and flying well.
Mon 17th May '10
Flying very well and working blossom in orchard and brassica flowers. Jim noticed bees in his garden on the Rosemary.
Sat 15th May '10
Very warm and sunny. Now have a thermometer T 15 C. Sprayed with homeopathy (6) in warmed rainwater. Only opened bars very small amount to do this. Removed feed as almost empty. Flying well and seem happy. Homeopathy now monthly for 3 months.
Weds 12th May '10
2.45 pm Joined by Essie and Richard (from course) to observe and help. Removed cross comb and attached to TB's. Removed one old frame and replaced with this TB. Also moved one of TB's with new comb on from last Weds to brood end as no activity on it. Removed more side bars from dadant frames. Left 2-3 on as bees were building comb in right direction to side of them. Dusted with icing sugar. Much less traumatic operation than last time. Bees less angry. Essie got one sting. Taken most of feed.
7 pm Farmer sprayed field at back of orchard with herbicides. Mick asked what it was - Hussar plus another one (didn't catch name will try to find out).
Sun 9th May '10
4.30 pm Some activity still despite drop in temperature. Powdered bees from underneath and above. New comb being built at R-angle to old comb on the frames again. Counted 12 dead bees and 2 at entrance that hadn't been removed.
Fresh comb that we moved had been patched up and extended slightly.
Fri 7th May '10
1.30 pm Put bait flowerpot out in case of swarm. Took old comb out of hive to put in this. Bees still very busy repairing their new comb that we moved. No feed taken as yet.
Thurs 6th May '10
Varroa count 50. This was after sugar dusting and moving fresh comb. Will do a 7 day count soon.
150 gms sugar in 150 ml rainwater given in contact feeder and pinch of salt. No chamomile this time.
Weds 5th May '10
Varroa count 10.
1 pm Weighed ( with roof on as forget to check what we did last time ) (R) 18 kgs (L) 19.5 Kgs
Replaced varroa counting board at 1 pm. Removed fresh comb that was at R-angle to dadant frames. This contained both drone and worker brood. Possible queen cell or more likely a queen cup. Varroa seen on pupae and in fresh comb cells. Sewed 2 bits onto TB's and put them back in hive.
Removed side bars from 3 frames. Will do rest at later date. Dusting with icing sugar carried out in brood area. Next dusting due on Sun 9th May.
Tues 4th May '10
6.40 pm Little activity as sun is low and temperature has dropped. Observed collection of 5 bees on the underside of varroa mesh. 4 look dead and 1 alive. Observed one dying bee repeatedly trying to clean it's abdomen with hind legs.
Mon 3rd May '10
1.30 pm Homeopathy (5) given. Some bees with small abdomens and dying! Also some with tatty wings. No sign of varroa. Bees building 3 combs at right angle to sidebars of the dadants, away from entrance, at stores end, across several frames! HELP!
Phoned Phil - Do varroa count for 24 hrs as may have K wing virus and varroa. Then do dusting with icing sugar over backs of bees in brood area. Then repeat 24 hr count. Will get board fixed underneath to do this. As for the 3 R-angled combs. These must be removed. Will try to sew to TB's and put back in brood area. Also try to take off second halves of dadant side bars. Check for drone and queen cells at same time.
7.30 pm Fitted floor invented by Mick to catch varroa. Fitted graph paper to it covered with Nature's World Petroleum free jelly (like Vaseline).
Mon 26th April '10
4 pm Rained yesterday, so didn't do treatment. Sprayed today instead using warmed rainwater and homeopathy (4).
Sat 24th April '10
4 pm Due for homeopathy today, but cooled down. Will leave until tomorrow.
Tues 20th April '10
3 - 4.30 pm Fitted varroa floor on second hive. Transferred bees into this. Saw queen in brood area. Bees very docile. One frame unattached from TB, so reattached. Lots of brood and stores. Some brace comb.
Set old hive up as possible trap for a swarm.
Mon 19th April '10
3 pm Very busy, crowded at entrance. Removed grass bung from 3rd hole so they can get in easier.
5.50 pm Little or no activity now. Cool northerly.
Sat 17th April ‘10
4 pm Sprayed with homeopathy (3). Lots of bees. Lots of activity. V sunny & warm. Homeopathy now weekly for 3 weeks.
Fri 16th April ‘10
12.15 pm Lots of activity. Many bees returning with pollen. Sunny, but with cold easterly wind going into entrance. Syrup feed only a few cms down.
5.30 pm Removed feed. Weighed without lid on: (R) 14kgs (L) 17kgs. Still flying well
Sprayed with homeopathy (2).
Thurs 15th April ‘10
10 am Not much of syrup gone – will leave in for now. Bees flying already. V sunny.
4.15 pm Sprayed with homeopathy (1). Peaked in, lots of activity. Grass stalks falling down in hive so removed. Mat seems to do the job on it's own.
Weds 14th April ‘10
Fed with sugar syrup in contact feeder.150g sugar / 150ml chamomile tea & pinch of salt. Very cold, no bees flying.
Checked late afternoon and they have started flying, as warmed up a bit.
Tues 13th April ‘10
Cleared a few more dead bees out. Opened follower board at stores end & looked in. Not started building new comb yet. Flying well.
Mon 12th April ‘10
Took roof off. 6 dead bees & debris on top of mesh. Removed these. Bees seem happy.
Sun 11th April ‘10
Last day of course. Bees escaping up between bars because wire ties causing gaps. Grass stalks packed between bars & mat put over the top.
Sat 10th April ‘10
Natural Beekeeping Course. Bees transferred to TBH by Phil. Bottom bars and ½ of side bars removed from frames. Trimmed at bottom to fit into TBH. Fixed to TB’s using wire. Very docile bees, aren't we lucky!
Thurs 8th April ’10
Nucleus delivered by Phil Chandler from Buckfast Abbey on ½ Dadant frames.